# MAPF-GPT: Imitation Learning for Multi-Agent Pathfinding at Scale

### v1

@@ Introduction
  • We present the largest MAPF dataset for decision-making, consisting of 1 billion observation-action pairs.
@@
Our approach, MAPF-GPT, is to learn to imitate an expert in solving MAPF. The learning phase of MAPF-GPT consists of the four major steps: creating MAPF scenarios, generating ground truth solutions, tokenizing these solutions, and executing the main training loop – see Fig. 1. We will now sequentially describe these steps.
@@
A large, curated dataset is crucial for any data-driven method including ours. To create the set of training instances we used POGEMA (Skrynnik et al. 2024a), a versatile tool for developing learnable MAPF solvers that includes utilities to generate maze-like maps and maps with random obstacles, as well as to create MAPF instances from them (i.e. assigning start-goal locations). For our purposes we generated 10K of maze-like maps and 2.5K random maps and further created 3.75M different problem instances on these maps. The size of the maps varied from $17\times 17$ to $21\times 21$, the number of the agents was $16$, $24$, or $32$. Please note that as we aim to create an individual policy to solve MAPF in a decentralized fashion (i.e. each agent makes its own decision on how to move based on its local observation) it is not the size of the maps that actually matters but rather the density, i.e. the ratio of the free space to the space occupied by the agents. We use moderate and considerably high densities to make the agents face challenging patterns requring coordination (especially on the mazes maps).
@@
The output of LaCAM on a single MAPF instance is the set of the individual plans. We, further, trace each plan and reconstruct an agent’s (local) observations in order to form (observation, action) pairs. We use the following post-processing to filter out some of them. First, if several pairs share the exactly the same observation we keep only one of them (picked randomly). Second, we observe that the fraction of the pairs when an agents waits at the goal is very high, because in many cases numerous agents wait for long times for other agents to reach their goals and just stand still. In fact, almost $40\%$ of the actions in the expert data were the waiting ones. To remove this imbalance we discard $80\%$ of wait-at-target actions. We end up with 900M of (observation, actions) pairs from the mazes maps, and 100M – from the random ones. A 9:1 proportion was chosen due to the maze maps possessing more challenging layouts with numerous narrow passages that require a high degree of cooperation between the agents.
We believe that the obtained dataset composed of 1B (observation, action) pairs is currently the largest dataset of such kind and may bring value to the other researchers developing learnable MAPF solvers. We are committed to open-source this dataset. Additional technical details are presented in the Appendix.
@@
The second part of the observation contains data about the agent itself and the nearby ones. The information about each agent consists of the coordinates of its current and goal locations, actions history, i.e., the actions that were made in the previous $k$ steps, and the action that is preferable w.r.t. the agent’s individual cost-to-go map – the so-called greedy action. Please note, that at maximum there can be 2 most preferable actions on a 4-connected grid. Thus we use a special marker to denote such “double” greedy actions (e.g. “up-right”).
@@
As the model backbone, we used a modern decoder-only transformer (Brown et al. 2020). We used a softmax layer to parameterize the discrete probability distribution. To sample an action, we then used multinomial sampling. The length of the input sequence (context size) of the model is $256$. The output size of the model is $5$ since each agent has $5$ available discrete actions. We used learnable position embeddings. Our largest model has $85$M parameters. We don’t use causal masking, which is common practice in NLP (Radford et al. 2019), since the model predicts only a single action ahead in a non-autoregressive manner. To speed up training, we used the flash attention technique (Dao 2024).
@@
The model was trained to clone the behavior of the expert policy using cross-entropy loss (i.e., log-loss) via mini-batch-based stochastic gradient descent, using AdamW (Loshchilov and Hutter 2019). The target label of such loss is a ground-truth action index provided by the expert policy LaCAM. LaCAM is a centralized approach that builds a path for all agents during the whole episode, providing information about the full environment state. In contrast, the trainable model uses the local observation $o$ of the agent $u$.
@@
We used a batch size of $512$ for the medium-sized model and $2048$ for the small-sized model. We trained the small-sized model for 30,000 iterations, which corresponds to 4 epochs on 150M data points. We used $2000$ warm-up iterations and cosine annealing (Loshchilov and Hutter 2017), with gradient clipping of $1.0$ and a weight decay parameter of $0.1$.
@@
In the first series of experiments, we compare 3 variants of MAPF-GPT varying in the number of parameters in their neural networks (2M, 6M, 85M) with the state-of-the-art learnable MAPF solvers: DCC (Ma, Luo, and Pan 2021) and SCRIMP (Wang et al. 2023) [^1] Please note that a plethora of learnable MAPF solvers exists however only few of them are tailored to the studied setting: MAPF with non-disappearing agents.. We use the pre-trained weights for DCC and SCRIMP. These weights were obtained by the authors while training on the random maps. Additionally, we present the results of LaCAM (Okumura 2024), which served as the expert centralized solver for data collection. For evaluation we used Random, Mazes, Warehouse, MovingAI maps (the latter two are out-of-distribution for all learnable solvers). The details on the maps and problem instances are given in Appendix.
@@
Figure 4 shows the Sum-of-Costs (SoC) achieved by the solvers relative to SoC of LaCAM (the lower - the better). As can be seen, MAPF-GPTs again outperform the other approaches and their performance correlates with the number of model parameters. Interestingly, on the Warehouse map, some of the instances are solved better than LaCAM did in terms of SoC.
@@
We evaluated MAPF-GPT in both zero-shot and fine-tuned configurations. To fine-tune the model, we generated an additional dataset using mazes maps. For expert data, we employed the RHCR approach (Li et al. 2021), as LaCAM is not well-suited for LMAPF. The dataset contains 90 million tensors. We used MAPF-GPT-6M for this experiment.
The results are presented in Table 2. Even the zero-shot model is able to compete with other existing learning-based approaches, such as Follower (Skrynnik et al. 2024b) and MATS-LP (Skrynnik et al. 2024c). Moreover, in all cases, fine-tuning improved the results of the MAPF-GPT-6M model, even outperforming other approaches on the warehouse map.
@@
MAPF-GPT-6M 1.4967 0.9081 2.3444 2.8404
MAPF-GPT-6M-t 1.5068 1.0874 2.7790 2.9938
RHCR 2.1636 1.5542 2.3522 3.4798
Follower 1.6370 1.1399 2.7306 3.2713
MATS-LP 1.6742 1.1253 1.7013 3.3202
@@ Conclusion
In this work we have studied MAPF problem as a sequential decision making problem. We proposed an approach to obtain a an individual policy based on state of the art in machine learning, i.e. (supervised) imitation learning on the expert data – MAPF-GPT. To train MAPF-GPT we create a comprehensive dataset of expert MAPF solutions, transformed these solutions into the (observation, action) pairs, tokenize them and learn several transformer (with the varying number of parameters) on this data. Empirically we show that even with a quite moderate number of parameters, i.e. 6M, MAPF-GPT notably outperforms modern learnable MAPF competitors across a large variety of setups. Our results provide a clear positive answer to the question “If it is possible to create a strong learnable MAPF solver purely by imitation learning”. The limitations of our approach are discussed in Appendix.
@@ Appendix B Implementation Details
The GPT model’s code is primarily based on the NanoGPT codebase [^2] https://github.com/karpathy/nanoGPT. NanoGPT was chosen for its straightforward and modular design, making it easier to modify and adapt the code, including potential adjustments for fine-tuning in environment (e.g. with reinforcement learning).
Appendix C Hyperparameters and Training Times
This section details the hyperparameters used while training all three our models: MAPF-GPT-2M, MAPF-GPT-6M, and MAPF-GPT-85M. Common parameters for the models are listed in Table 3, while Table 4 provides model-specific hyperparameters for MAPF-GPT-2M, MAPF-GPT-6M, and MAPF-GPT-85M.
Table 3: Common Hyperparameters for MAPF-GPT Models
@@ Appendix C Hyperparameters and Training Times
Table 4: Model-Specific Hyperparameters
@@
Number of epochs 46.875 12.5 5.625
Total training iterations 15,000 30,000 360,000
@@
We didn’t change the hyperparameters between the models except for those that influence their size and the duration of the learning process. Training the 85M model for 360K iterations took 85 hours using 4x H100 80GB NVIDIA GPUs. Training the 6M model for 30K iterations took 50 hours using 2x A100 80GB NVIDIA GPUs. The 2M model was trained on a single H100 80GB NVIDIA GPU within 12 hours. It is also worth noting that while training the 6M and 2M models we could not fully utilize the computational power of the corresponding GPUs, as the bottleneck was actually in data processing and transferring it to the GPU.
(a) Mazes
(b) Random
(c) Puzzles
(d) MovingAI
(e) Warehouse
Figure 6: Examples of maps presented in the POGEMA Benchmark.
@@ Appendix D Dataset: Technical Details
The resulting log files, which contain the actions taken by the agents at each time step, were processed to generate local observations for every agent. To speed up the generation process, we processed the log files for several problem instances in parallel. At this stage, we removed redundant waiting actions and duplicate observations as detailed in the main body of the paper. The resultant data was saved in .hdf5 format with compression to save disk space.
In the final step, 12,500 .hdf5 files were shuffled and split into $50$ chunks. The data from these chunks was also shuffled and saved into $10$arrow files per chunk. Each .arrow file contains $2^{21}$(slightly more than 2 million) observation-action pairs, with 10% of the data coming from the random maps and 90% – from mazes.
The resultant 1B dataset contains 500 such files, requiring 258GB of disk space. While training the MAPF-GPT-6M and MAPF-GPT-2M models, we randomly selected 75 and 20 files from this dataset, resulting in 150M and 40M datasets, respectively. The sizes of the datasets were selected according to Chinchilla paper (Hoffmann et al. 2022), which found scaling low of 20 tokens per model parameter, showing empirical superiority of model of 70 billion tokens, trained on 1.4 trillion tokens. This research found that using about 20 tokens for each parameter in a language model strikes the best balance between model size and the amount of training data, leading to better performance and more efficient training. The resultant 1B dataset contains 500 such files, requiring 258GB of disk space. While training the MAPF-GPT-6M and MAPF-GPT-2M models, we randomly selected 75 and 20 files from this dataset, resulting in 150M and 40M datasets, respectively. The full dataset was used for training MAPF-GPT-85M model. The sizes of the datasets were selected according to Chinchilla paper (Hoffmann et al. 2022), which empirically found a scaling law of 20 tokens per model parameter, demonstrating the efficiency of a 70-billion-parameter model trained on 1.4 trillion tokens compared to larger models trained on fewer tokens.
@@ Appendix E Benchmark
For empirical evaluation we have utilized the evaluation tools provided by the POGEMA benchmark(Skrynnik et al. 2024a). Specifically, we used the following set of maps: Random, Mazes, Warehouse (Li et al. 2021), MovingAI and Puzzles. Table 5 contains information about the number of agents, maps, and instances per map (seeds) used in each set.
@@
Agents Maps MapSize Seeds MaxSteps
Random [8, 16, 24, 32, 48, 64] 128 17$\times$17 - 21$\times$21 1 128
Mazes [8, 16, 24, 32, 48, 64] 128 17$\times$17 - 21$\times$21 1 128
Warehouse [32, 64, 96, 128, 160, 192] 1 33$\times$46 128 128
MovingAI [64, 128, 192, 256] 128 64$\times$64 1 256
Puzzles [2, 3, 4] 16 5$\times$5 10 128
Figure 6 showcases examples of maps from the POGEMA benchmark. The Mazes and Random map sets are generated using the built-in generators within POGEMA. The Puzzles maps are handcrafted. The Warehouse is a single map with limitations on possible start and goal positions, which has been used in previous LMAPF-related studies (Li et al. 2021; Skrynnik et al. 2024b). Start locations on this map can only be generated on the left or right sides of the map, where there are no obstacles, while goal locations can only be placed beyond or below the obstacles in the center of the map. These constraints prevent the generation of instances with more than 192 agents on this map. The final evaluated set of maps is MovingAI, referred to as MovingAI-tiles in the POGEMA benchmark. While the original MovingAI dataset includes various types of maps, including random and maze maps, the MovingAI-tiles set in POGEMA contains only city maps. Additionally, due to the large size ($256\times 256$) of city maps, they were divided into 16 tiles, resulting in maps of $64\times 64$ size.
@@
Appendix F Evaluation on Puzzles

### v2

@@ Introduction
  • We present the largest MAPF dataset for decision-making, consisting of 1 billion (observation, action) pairs.
@@
Our approach, MAPF-GPT, is to learn to imitate an expert in solving MAPF. The learning phase of MAPF-GPT consists of the four major steps: creating MAPF scenarios, generating ground truth solutions, tokenizing these solutions, and executing the main training loop – see Figure 1. We will now sequentially describe these steps.
@@
A large, curated dataset is crucial for any data-driven method including ours. To create the set of training instances we used POGEMA (Skrynnik et al. 2024a), a versatile tool for developing learnable MAPF solvers that includes utilities to generate maze-like maps and maps with random obstacles, as well as to create MAPF instances from them (i.e. assigning start-goal locations). For our purposes we generated 10K of maze-like maps and 2.5K random maps and further created 3.75M different problem instances on these maps. The size of the maps varied from $17\times 17$ to $21\times 21$, the number of the agents was $16$, $24$, or $32$. Please note that as we aim to create an individual policy to solve MAPF in a decentralized fashion (i.e. each agent makes its own decision on how to move based on its local observation) it is not the size of the maps that actually matters but rather the density, i.e. the ratio of the free space to the space occupied by the agents. We use moderate and considerably high densities to make the agents face challenging patterns requiring coordination (especially on the maze-like maps).
@@
The output of LaCAM on a single MAPF instance is the set of the individual plans. We, further, trace each plan and reconstruct an agent’s (local) observations in order to form (observation, action) pairs. We use the following post-processing to filter out some of them. First, if several pairs share the exactly the same observation we keep only one of them (picked randomly). Second, we observe that the fraction of the pairs when an agents waits at the goal is very high, because in many cases numerous agents wait for long times for other agents to reach their goals and just stand still. In fact, almost $40\%$ of the actions in the expert data were the waiting ones. To remove this imbalance we discard $80\%$ of wait-at-target actions. We end up with 900M of (observation, action) pairs from the maze-like maps, and 100M – from the random ones. A 9:1 proportion was chosen due to the maze maps possessing more challenging layouts with numerous narrow passages that require a high degree of cooperation between the agents.
We believe that the obtained dataset composed of 1B (observation, action) pairs is currently the largest dataset of such kind and may bring value to the other researchers developing learnable MAPF solvers. Additional technical details are presented in the Appendix.
@@
The second part of the observation contains data about the agent itself and the nearby ones. The information about each agent consists of the coordinates of its current and goal locations, actions history, i.e., the actions that were made in the previous $k$ steps, and the action that is preferable w.r.t. the agent’s individual cost-to-go map – the so-called greedy action. Please note that there may be cases where more than one action leads to a decrease in cost-to-go. Thus, we use special markers to indicate these ”multi-direction” greedy actions (e.g., ”up-right”).
@@
As the model backbone, we used a modern decoder-only transformer (Brown et al. 2020). We used a softmax layer to parameterize the discrete probability distribution. To sample an action, we then used multinomial sampling. The length of the input sequence (context size) of the model is $256$. The output size of the model is $5$ since each agent has $5$ available discrete actions. We used learnable position embeddings. Our largest model has $85$M parameters. Besides it, we also trained models with $6$M and $2$M parameters. We don’t use causal masking, which is common practice in NLP (Radford et al. 2019), since the model predicts only a single action ahead in a non-autoregressive manner. To speed up training, we used the flash attention technique (Dao 2024).
@@
The model was trained to replicate the behavior of the expert policy using cross-entropy loss (i.e., log-loss) via mini-batch stochastic gradient descent, optimized with AdamW (Loshchilov and Hutter 2019). The target label for this loss is the ground-truth action index provided by the expert policy, LaCAM. LaCAM is a centralized solver that builds a path for all agents during the whole episode, leveraging information about the full environment state. In contrast, the trainable model relies solely on the local observation $o$ of each agent $u$.
@@
We used $2000$ warm-up iterations and cosine annealing (Loshchilov and Hutter 2017), with a gradient clipping value of $1.0$ and a weight decay parameter of $0.1$. The entire 1B dataset was utilized to train the 85M model, which underwent 1M iterations with a batch size of $512$, resulting in $15.625$ epochs based on the gradient accumulation steps, set at $16$. For training the 6M and 2M models, we used portions of the 1B dataset – 150M and 40M, respectively. Additional details about the parameters influencing the training process are provided in the Appendix.
@@
In the first series of experiments, we compare 3 variants of MAPF-GPT varying in the number of parameters in their neural networks (2M, 6M, 85M) with the state-of-the-art learnable MAPF solvers: DCC (Ma, Luo, and Pan 2021) and SCRIMP (Wang et al. 2023) [^1] Please note that a plethora of learnable MAPF solvers exists however only few of them are tailored to the studied setting: MAPF with non-disappearing agents.. We use the pre-trained weights for DCC and SCRIMP. These weights were obtained by the authors while training on the random maps. Additionally, we present the results of LaCAM (Okumura 2024), which served as the expert centralized solver for data collection. For evaluation we used Random, Mazes, Warehouse, MovingAI maps (the latter two are out-of-distribution for all learnable solvers). The details on the maps and problem instances are given in the Appendix.
@@
Figure 4 shows the Sum-of-Costs (SoC) achieved by the solvers relative to SoC of LaCAM (the lower - the better). As can be seen, MAPF-GPTs again outperform the other approaches and their performance correlates with the number of model parameters. Interestingly, there are some rare cases on Random maps where DCC and MAPF-GPT-85M outperformed LaCAM in terms of SoC. The same situation is observed for MAPF-GPT-6M on Mazes maps.
@@
We evaluated MAPF-GPT in both zero-shot and fine-tuned configurations. To fine-tune the model, we generated an additional dataset using Mazes maps. For expert data, we employed the RHCR approach (Li et al. 2021), as LaCAM is not well-suited for LMAPF. The dataset contains 90 million (observation, action) pairs. We used MAPF-GPT-6M for this experiment.
The results are presented in Table 2. Even the zero-shot model is able to compete with other existing learning-based approaches, such as Follower (Skrynnik et al. 2024b) and MATS-LP (Skrynnik et al. 2024c). Moreover, in all cases, fine-tuning improved the results of the MAPF-GPT-6M model.
@@
MAPF-GPT-6M 1.497 0.908 1.113 2.840
MAPF-GPT-6M-t 1.507 1.087 1.270 2.994
RHCR 2.164 1.554 2.352 3.480
Follower 1.637 1.140 2.731 3.271
MATS-LP 1.674 1.125 1.701 3.320
@@ Conclusion
In this work, we studied the MAPF problem as a sequential decision-making task. We proposed an approach to derive an individual policy based on state-of-the-art machine learning techniques, specifically (supervised) imitation learning from expert data —- MAPF-GPT. To train MAPF-GPT, we created a comprehensive dataset of expert MAPF solutions, transformed these solutions into (observation, action) pairs, tokenized them, and trained several transformer models (with varying numbers of parameters) on this data. Empirically, we demonstrate that even with a quite moderate number of parameters, such as 2M, MAPF-GPT significantly outperforms modern learnable MAPF competitors across a wide range of setups. Our results provide a clear positive answer to the question, ”Is it possible to create a strong learnable MAPF solver purely through imitation learning?”. The limitations of our approach are discussed in the Appendix.
@@ Appendix B Implementation Details
The GPT model’s code is primarily based on the NanoGPT codebase [^2] https://github.com/karpathy/nanoGPT. NanoGPT was chosen for its simple yet modular design, making it easier to modify and adapt the code, including potential adjustments for fine-tuning in environment (e.g., with reinforcement learning).
Appendix C Hyperparameters and Training Details
This section details the hyperparameters used while training all three our models: MAPF-GPT-2M, MAPF-GPT-6M, and MAPF-GPT-85M. Common parameters for the models are listed in Table 3.
Table 3: Common hyperparameters for MAPF-GPT models.
@@ Appendix C Hyperparameters and Training Details
Table 4 provides model-specific hyperparameters for MAPF-GPT-2M, MAPF-GPT-6M, and MAPF-GPT-85M. We didn’t change the hyperparameters between the models except for those that influence their size and the duration of the learning process. Training the 85M model for 1M iterations took 243 hours using 4x H100 80GB NVIDIA GPUs. Training the 6M model for 30K iterations took 50 hours using 2x A100 80GB NVIDIA GPUs. The 2M model was trained on a single H100 80GB NVIDIA GPU within 12 hours. It is also worth noting that while training the 6M and 2M models we could not fully utilize the computational power of the corresponding GPUs, as the bottleneck was actually in data processing and transferring it to the GPU.
Table 4: Model-specific hyperparameters.
@@
Number of epochs 46.875 12.5 15.625
Total training iterations 15,000 30,000 1,000,000
@@
The loss curves for MAPF-GPT-85M are presented in Figure 6, showing the averaged loss for both training and validation. The validation dataset contains an equal 1:1 ratio of (observation, action) pairs between maze-like and random maps, in contrast to the 9:1 ratio in the training dataset.
Figure 6: Training and validation loss curves for MAPF-GPT-85M.
During training, we also performed evaluations in the environment using intermediate checkpoints. The aggregated results of these evaluations are presented in Figure 7. Intermediate evaluations were conducted on 6 random maps with 48 and 64 agents, as well as 6 maze maps with 32 and 48 agents. These maps were taken from the set used to generate the validation dataset.
Figure 7: Aggregated evaluation results for Success rate and Sum of Costs during training.
@@ Appendix D Dataset: Technical Details
The resulting log files, which contain the actions taken by the agents at each time step, were evenly distributed into $50$ chunks. The logs from each chunk were processed to generate local observations for every agent. At this stage, we removed redundant waiting actions and duplicate observations as detailed in the main body of the paper.
In the final step, the obtained data was shuffled and saved into $10$arrow files per chunk. Each .arrow file contains $2^{21}$(slightly more than 2 million) (observation, action) pairs, with 10% of the data coming from the random maps and 90% – from mazes.
The resultant 1B dataset contains 500 such files, requiring 258GB of disk space. While training the MAPF-GPT-6M and MAPF-GPT-2M models, we randomly selected 75 and 20 files from this dataset, resulting in 150M and 40M datasets, respectively. The full dataset was used for training MAPF-GPT-85M model. The sizes of the datasets were selected according to Chinchilla paper (Hoffmann et al. 2022), which empirically found a scaling law of 20 tokens per model parameter, demonstrating the efficiency of a 70-billion-parameter model trained on 1.4 trillion tokens compared to larger models trained on fewer tokens.
@@ Appendix E Benchmark
For empirical evaluation we have utilized the evaluation tools provided by the POGEMA benchmark (Skrynnik et al. 2024a). Specifically, we used the following set of maps: Random, Mazes, Warehouse (Li et al. 2021), MovingAI (Stern et al. 2019) and Puzzles. Table 5 contains information about the number of agents, maps, and instances per map (seeds) used in each set.
@@
Set Agents Maps Map Size Seeds Steps
Random 8, 16, 24, 32, 48, 64 128 17$\times$17 - 21$\times$21 1 128
Mazes 8, 16, 24, 32, 48, 64 128 17$\times$17 - 21$\times$21 1 128
Warehouse 32, 64, 96, 128, 160, 192 1 33$\times$46 128 128
MovingAI 64, 128, 192, 256 128 64$\times$64 1 256
Puzzles 2, 3, 4 16 5$\times$5 10 128
(a) Random
(b) Mazes
(c) Warehouse
Figure 8: The examples of maps from the POGEMA benchmark. Plots show the maximum number of agents tested on each map: 64 agents for Random and Mazes sets, and 192 agents for Warehouse scenario.
Figure 8 showcases examples of maps from the POGEMA benchmark. The Mazes and Random map sets are generated using the built-in generators within POGEMA. The Puzzles maps are handcrafted. The Warehouse is a single map with limitations on possible start and goal positions, which has been used in previous LMAPF-related studies (Li et al. 2021; Skrynnik et al. 2024b). Start locations on this map can only be generated on the left or right sides of the map, where there are no obstacles, while goal locations can only be placed beyond or below the obstacles in the center of the map. These constraints prevent the generation of instances with more than 192 agents on this map.
The final evaluated set of maps is MovingAI, referred to as MovingAI-tiles in the POGEMA benchmark. While the original MovingAI dataset includes various types of maps, including random and maze maps, the MovingAI-tiles set in POGEMA contains only city maps. Additionally, due to the large size ($256\times 256$) of city maps, they were divided into 16 tiles, resulting in maps of $64\times 64$ size. An example of such scenario is presented in Figure 9.
Figure 9: Example of a city map from the MovingAI-tiles set in the POGEMA benchmark. The original city maps, with dimensions of $256\times 256$, were divided into 16 tiles, resulting in smaller $64\times 64$ maps.
@@
Appendix F Limitations
The main limitation of MAPF-GPT is a generic one, shared with the other learnable methods, i.e. it lacks theoretical guarantees. The next limitation is that training large models, e.g., MAPF-GPT-85M, is quite demanding (requires expensive hardware and prolonged time). Still, our smaller models containing 6M and 2M parameters are much less demanding while providing competitive results. It should also be noted that all MAPF-GPT models are sensitive to the quality of trajectories in the expert data set. As has been shown in previous research on behavior cloning with transformers, e.g. (Chen et al. 2021), adding low-quality trajectories to expert data may lead to significant degradation in model’s performance. It is also unclear how effectively MAPF-GPT can replicate the behavior of the other existing centralized approaches (such as CBS (Sharon et al. 2015) that is an optimal MAPF solver). This dependence on the type of behavioral expert policy requires further research.
Appendix G Evaluation on Puzzles